### v4

# Minimal codes from hypersurfaces in even characteristic

Abstract
The setting of projective systems can be used to study the parameters of a projective linear code $\mathscr{C}$. This can be done by considering the intersections of the point set $\Omega$ defined by the columns of a generating matrix for $\mathscr{C}$ with the hyperplanes of a projective space. In particular, $\mathscr{C}$ is minimal if $\Omega$ is cutting set, i.e., every hyperplane is spanned by its intersection with $\Omega$. Minimal linear codes have important applications for secret sharing schemes and secure two-party computation.
In this article we first investigate the properties of some algebraic hypersurfaces $\mathscr{V}_{\varepsilon}^{r}$ related to certain quasi-Hermitian varieties of $\mathrm{PG}(r,q^{2})$, with $q=2^{e}$, $e>1$ odd. These varieties give rise to a new infinite family of linear codes which are minimal except for $r=3$ and $e\equiv 1\pmod{4}$.
In the case $r\in\{3,4\}$, we exhibit codes having at most 6 non-zero weights whose we provide the complete list. As a byproduct, we obtain $(r+1)$-dimensional codes with just $3$ non-zero weights. We point out that linear codes with few weights are also important in authentication codes and association schemes.
In the last part of the paper, we consider an extension of the notion of being cutting with respect to subspaces other than hyperplanes and introduce the definition of a cutting gap in order to characterize and measure what happens when this property is not satisfied. Finally, we apply these concepts to both Hermitian varieties and the hypersurfaces $\mathscr{V}_{\varepsilon}^{r}$.
Keywords: Algebraic variety, Quasi-Hermitian variety, linear code.
MSC Classification: 94B05, 51A05, 51E20.
1 Introduction
Linear codes derived from projective geometry have been extensively investigated due to the rich interplay between algebraic and geometric structures. The approach of constructing generator matrices for codes by considering the coordinates of point-sets in projective spaces has been pioneered by [19] and [8] and it has proved to be extremely fruitful ever since. In particular, given a non-degenerate $q$-ary linear $[n,r,d]$-code $\mathscr{C}$, the projective system associated with $\mathscr{C}$ is defined as the (multi)set $\Omega$ of points in $\mathrm{PG}(r-1,q)$ corresponding to the columns of a generator matrix of $\mathscr{C}$. Conversely, any such multiset $\Omega$ yields a code $\mathscr{C}(\Omega)$ by taking the coordinates of the points in $\Omega$ as the columns of a generator matrix. The resulting code $\mathscr{C}(\Omega)$ is unique up to monomial equivalence and is referred to as the code generated or induced by $\Omega$.
This geometric approach is particularly effective when $\Omega$ arises from algebraic or geometric structures, as it often leads to codes with desirable features, such as minimality or a small number of distinct weights relative to the code length.
A linear $[n,r,d]$-code is said to be minimal if every non-zero codeword is minimal, i.e., uniquely determined (up to a non-zero scalar multiple) by its support. Geometrically, this corresponds to the condition that the associated projective system $\Omega$ is a cutting set, meaning that every hyperplane of $\mathrm{PG}(r-1,q)$ is spanned by its intersection with $\Omega$. Minimal codes are of interest not only for their geometric significance but also for their applications in cryptography, including secret-sharing schemes and secure two-party computation, see [20, 12].
In this paper, we construct new families of linear codes with few weights, derived from hypersurfaces associated with quasi-Hermitian varieties in even characteristic. These constructions are noteworthy for their geometric novelty. Specifically, the hypersurfaces considered are linked to Buekenhout–Tits quasi-Hermitian varieties (or BT quasi-Hermitian varieties, for short), which exist in every $\mathrm{PG}(r-1,q^{2})$ with $r>2$ and $q=2^{e}$, where $e\geq 3$ is odd. The structure of these varieties enables the exploration of new geometric configurations in finite projective spaces.
In certain cases, the associated codes are minimal; in others, they exhibit exactly three distinct weights. Codes with few weights are particularly valuable in the design of authentication codes and association schemes (see [9]), and their weight distribution plays a crucial role in evaluating decoding performance and efficiency.
In the final part of the paper, we introduce a new concept, the cutting gap, to quantify the extent to which a set $\Omega$ fails to satisfy the cutting property (i.e. that a subspace is generated by its intersection with $\Omega$). This notion aims to quantify how much a given projective system $\Omega$ deviates from being a strong blocking set (see [13]) and also aims to open the way to extending the concept of minimality to higher Hamming weights.
The paper is organized as follows. In Section 2, we recall foundational results on projective codes and quasi-Hermitian varieties, with particular emphasis on BT quasi-Hermitian varieties. Section 3 is devoted to the analysis of hypersurfaces $\mathscr{V}^{r}_{\varepsilon}$ associated with these varieties, and we determine the conditions under which they yield minimal codes. For $r=3$, we compute the nonzero weights of $\mathscr{C}_{\varepsilon}^{3}:=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$, while for $r\geq 3$, we construct a new family of three-weight codes.
In Section 4, we formally define the concept of cutting gap and explore its implications. We apply this framework to Hermitian varieties and study the intersections of $\mathscr{V}^{4}_{\varepsilon}$ with planes and lines in $\mathrm{PG}(4,q^{2})$, including an analysis of Fermat curves. As a result, we provide the complete sequence of cutting gaps for $\mathscr{V}^{4}_{\varepsilon}$ and the full list of weights for the code $\mathscr{C}_{\varepsilon}^{4}:=\mathscr{C}(\mathscr{V}^{4}_{\varepsilon})$. In both cases, $r=3$ or $r=4$, the limit
$\lim_{q\to\infty}\frac{N(q)}{d(q)}=1,$
where $N(q)$ denotes the code length and $d(q)$ the minimum distance. This implies that the asymptotic minimum distance of the constructed codes is relatively good. In the following table, we summarize all the results obtained concerning the code $\mathscr{C}_{\varepsilon}^{r}$.
Code Result
  $\mathscr{C}^{3}_{\varepsilon}$ Theorem 3.2: length
Theorem 3.3: weights for $e\equiv 3\pmod{4}$
Theorem 3.4: weights for $e\equiv 1\ \pmod{4}$
Theorem 3.5: minimality for $e\equiv 3\pmod{4}$
Corollary 3.6: $m$-th higher minimum weights
$\mathscr{C}^{r}_{\varepsilon},r\geq 4$ Theorem 3.12 : minimality for $e$ odd
$\mathscr{C}(\overline{\mathscr{V}_{\varepsilon}^{r}})$ Theorem 3.13: weights
$\mathscr{C}^{4}_{\varepsilon}$ Proposition 4.9: minimality, length and weights
2 Preliminaries
All codes studied in this paper are defined in the Hamming metric. In order to make the article more readable, we shall briefly recall in this section some classical notions from coding theory. For more details, the reader can refer to [18, 23]. We shall also recall here the construction of BT quasi-Hermitian varieties, to be used in the remainder of the work.
2.1 Coding Theory
Let $\mathscr{C}$ be a $q$-ary linear $[n,r,d]$-code, i.e. a $r$-dimensional linear $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance $d$. The support of a codeword $\boldsymbol{c}=(c_{1},\dots,c_{n})\in\mathscr{C}$ is the set $\mathrm{supp}(\boldsymbol{c}):=\{i:c_{i}\neq 0\}$. The support $\mathrm{supp}(\mathscr{C})$ of a code $\mathscr{C}$ is the union of the supports of its codewords. The code $\mathscr{C}$ is called non-degenerate if $\mathrm{supp}(\mathscr{C})=\{1,\ldots,n\}$; degenerate otherwise. The dual of an $[n,r,d]$-linear code $\mathscr{C}$ is defined as
$\mathscr{C}^{\perp}=\{\boldsymbol{c}^{\prime}\in\mathbb{F}_{q}^{n}\colon\boldsymbol{c}\cdot\boldsymbol{c}^{\prime}=0\quad\forall\boldsymbol{c}\in\mathscr{C}\},$
where $\cdot$ is the standard dot product of $\mathbb{F}_{q}^{n}$. A code is non-degenerate if its dual has minimum distance at least $2$.
Definition 2.1.
A codeword $\boldsymbol{c}\in\mathscr{C}$ is minimal if for any non-zero codeword $\boldsymbol{c^{\prime}}\in\mathscr{C}$ with $\mathrm{supp}(\boldsymbol{c^{\prime}})\subseteq\mathrm{supp}({\boldsymbol{c}})$ we have $\boldsymbol{c^{\prime}}\in\langle\boldsymbol{c}\rangle$. A code $\mathscr{C}$ is minimal if all of its non-zero codewords are minimal.
A code $\mathscr{C}$ is projective if its dual has minimum distance at least $3$. We observe that a $q$-ary $[n,r,d]$ code $\mathscr{C}$ is projective if and only if the columns of any of its generator matrices are pairwise non-proportional, see [10]. In particular, they can be regarded as the coordinates of a set $\Omega$ of $n$ distinct points in $\mathrm{PG}(r-1,q)$. We shall call $\Omega$ the projective system of $\mathscr{C}$. Conversely, given a set $\Omega$ of $n$ distinct points in $\mathrm{PG}(r-1,q)$ we can define a linear $[n,r]$-code $\mathscr{C}:=\mathscr{C}(\Omega)$ whose generator matrix has as columns the coordinates of the points of $\Omega$. Clearly, $\mathscr{C}(\Omega)$ is not uniquely defined as a code, but it is none-the-less unique up to monomial equivalence.
We recall the following result from from [4] and [22]:
Theorem 2.2.
Let $\Omega$ be a set of $N$ points of $\mathrm{PG}(r-1,q)$ such that $\langle\Omega\rangle=\mathrm{PG}(r-1,q)$. For each $i\in\{1,\ldots,N\}$, let $\boldsymbol{\textnormal{v}}_{i}\in\mathbb{F}_{q}^{r}$, be a fixed representative of a point $P_{i}=\langle\textbf{{v}}_{i}\rangle\in\Omega$ and denote by $\mathscr{C}(\Omega)$ the projective linear code having generator matrix whose columns are the vectors $\textbf{{v}}_{i}$. Then $\mathscr{C}(\Omega)$ is a minimal code if and only if for any hyperplane $\Pi$ of $\mathrm{PG}(r-1,q)$,
$\langle\Pi\cap\Omega\rangle=\Pi,$
that is $\Omega$ is cutting.
Ashikhmin and Barg in [5] provided a sufficient condition so that an $[n,r]$-linear code $\mathscr{C}$ over $\mathbb{F}_{q}$ is minimal, namely
$\frac{w_{\min}}{w_{\max}}>\frac{q-1}{q},$ (1)
where $w_{\min}$ and $w_{\max}$ are respectively the minimum and maximum weight of non-zero codewords of $\mathscr{C}$.
The intersection numbers of a projective system with the hyperplanes is closely related to the Hamming weights of a code. In particular the minimum distance of a code $\mathscr{C}=\mathscr{C}(\Omega)$, with $\Omega$ a projective system, is
$d=|\Omega|-\max\{|\Omega\cap\Pi|:\Pi\text{ hyperplane of }\mathrm{PG}(\langle\Omega\rangle)\}.$
Generalized Hamming weights for linear codes have been introduced by Wei in [24] as a way to characterize the performance for a particular type of a linear code.
Definition 2.3.
Let $\mathscr{C}$ be a code and $\mathscr{D}$} a subcode of $\mathscr{C}$. The $k$-th generalized Hamming weight of $\mathscr{C}$ is
$d_{k}(\mathscr{C}):=\min\{|\mathrm{supp}({\mathscr{D}})|\colon{\mathscr{D}}\leq\mathscr{C},\dim(\mathscr{D})=k\}.$
Clearly, $d_{1}(\mathscr{C})$ corresponds to the usual minimum Hamming distance $d$ of $\mathscr{C}$. More in general, it can be seen that for $1\leq k\leq r-1$ the kk–th Hamming weight of a projective code $\mathscr{C}=\mathscr{C}(\Omega)$ is given by
$d_{k}(\mathscr{C})=|\Omega|-\max\{|\Omega\cap\Pi|:\mathrm{codim}(\Pi)=k\}.$
2.2 Quasi-Hermitian varieties
Quasi-Hermitian varieties in $\mathrm{PG}(r,q^{2})$ are point sets having the same intersection numbers with respect to hyperplanes as a non-singular Hermitian variety $\mathscr{H}(r,q^{2})$ of $\mathrm{PG}(r,q^{2})$. Notice that a point set $\mathscr{S}$ in $\mathrm{PG}(r,q^{2})$, with $r>2$, having the same intersection numbers with respect to hyperplanes as a non-singular Hermitian variety also has the same number of points as $\mathscr{H}(r,q^{2})$ (see [21]).
These varieties can be seen as a natural extension to higher dimensions of the notion of a unital in the projective plane.
A unital embedded in $\mathrm{PG}(2,q^{2})$ is a set of $q^{3}+1$ points such that each line intersects it in either $1$ or $q+1$ points. The Hermitian curve $\mathscr{H}(2,q^{2})$ is a unital, known as the classical unital.
More precisely, the cardinality of a quasi-Hermitian variety in $\mathrm{PG}(r,q^{2})$ is
$|\mathscr{H}(r,q^{2})|=\frac{(q^{r+1}+(-1)^{r})(q^{r}-(-1)^{r})}{q^{2}-1},$
while its intersection numbers with respect to hyperplanes are respectively
$|\mathscr{H}(r-1,q^{2})|=\frac{(q^{r}+(-1)^{r-1})(q^{r-1}-(-1)^{r-1})}{q^{2}-1},$
and
$|\Pi_{0}\cap\mathscr{H}(r-2,q^{2})|=\frac{(q^{r}+(-1)^{r-1})(q^{r-1}-(-1)^{r-1})}{q^{2}-1}+(-1)^{r-1}q^{r-1},$
where $\Pi_{0}$ denotes a hyperplane tangent to the variety.
A non-singular Hermitian variety of $\mathrm{PG}(r,q^{2})$ is, by definition, a quasi-Hermitian variety: the classical quasi-Hermitian variety.
Angela Aguglia introduced in [1] a new family of non-classical quasi-Hermitian varieties which are linked to Buekenhout-Tits unitals. We shall call these varieties BT quasi hermitian varieties. They can be constructed as follows.
Let $e$ be an odd number, $r\geq 2$ and consider the projective space $\mathrm{PG}(r,q^{2})$ of order $q^{2}=2^{2e}$ and dimension $r$. Denote by
$(X_{0},X_{1},\ldots,X_{r}),$
the homogeneous coordinates of its points. Fix the hyperplane
$\Pi_{\infty}:=\{(X_{0},X_{1},\ldots,X_{r})\in\mathrm{PG}(r,q^{2})|\,X_{0}=0\}$
as the hyperplane of the ‘at infinity’ and denote by $\mathrm{AG}(r,q^{2})$ the affine space $\mathrm{PG}(r,q^{2})\setminus\Pi_{\infty}$ whose points have coordinates $(x_{1},\ldots,x_{r})$ where $x_{i}=X_{i}/X_{0}$ for $i=1,\ldots,r$.
Write
$\mathrm{Tr}_{q^{2}/q}(x):=x+x^{q},\qquad\mathrm{N}_{q^{2}/q}(x):=x^{q+1}\qquad\text{ for all }x\in\mathbb{F}_{q^{2}},$
for respectively the trace and norm functions of $\mathbb{F}_{q^{2}}$ over $\mathbb{F}_{q}$.
Choose now $\varepsilon\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ such that there exists a $\delta\in\mathbb{F}_{q}\setminus\{1\}$, $\mathrm{Tr}_{q/2}(\delta)=1$ that satisfies $\varepsilon^{2}+\varepsilon+\delta=0$. Then, it is straightforward to see that $\mathrm{Tr}_{q^{2}/q}(\varepsilon)=1$.
If $q=2^{e}$ with $e$ an odd number, then the function $\sigma:x\rightarrow x^{2^{\frac{e+1}{2}}}$ is an automorphism of $\mathbb{F}_{q}$.
Let now $\mathscr{V}^{r}_{\varepsilon}$ be the projective variety of $\mathrm{PG}(r,q^{2})$ with affine equation
$x_{r}^{q}+x_{r}=\Gamma_{\varepsilon}(x_{1})+\ldots+\Gamma_{\varepsilon}(x_{r-1})$ (2)
where
$\Gamma_{\varepsilon}(x)=[x+(x^{q}+x)\varepsilon]^{\sigma+2}+(x^{q}+x)^{\sigma}+(x^{2q}+x^{2})\varepsilon+x^{q+1}+x^{2}.$
Let also $\mathscr{H}^{r}_{\varepsilon}$ be the variety whose affine points are the same as those of $\mathscr{V}^{r}_{\varepsilon}$, but whose points at infinity are given by the Hermitian cone
$\mathscr{H}^{r}_{\varepsilon,\infty}:=\{(0,X_{1},\ldots,X_{r})|\quad X_{1}^{q+1}+\ldots+X_{r-1}^{q+1}=0\}.$ (3)
It is shown in [1] that $\mathscr{H}^{r}_{\varepsilon}$ is a non-classical quasi-Hermitian variety of $\mathrm{PG}(r,q^{2})$ which, for $r=2$, is a Buekenhout-Tits unital. We call $\mathscr{H}^{r}_{\varepsilon}$ a Buekenhout-Tits or BT quasi-Hermitian variety of $\mathrm{PG}(r,q^{2})$ for $r\geq 2$.
Note that, since $\mathscr{H}^{r}_{\varepsilon}$ is a quasi-Hermitian variety and it has the same affine points of $\mathscr{V}^{r}_{\varepsilon}$, every hyperplane $\pi\neq\Pi_{\infty}$ in $\mathrm{PG}(r,q^{2})$ meets $\mathscr{V}^{r}_{\varepsilon}$ in at least one affine point.
Theorem 2.4.
The number of affine points of $\mathscr{V}^{r}_{\varepsilon}$ is $q^{2r-1}$.
Proof.
It follows trivially from the fact that $\mathscr{H}_{\varepsilon}$ is a quasi-Hermitian variety, $\mathscr{H}^{r}_{\varepsilon,\infty}$ is a cone over $\mathscr{H}(r-2,q^{2})$ and $|\mathscr{V}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})|=|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})|$. ∎
2.3 Bounds on the number of points of hypersurfaces
Let $\theta_{q}(r)$ be the number of points of the $r$-dimensional finite projective space $\mathrm{PG}(r,q)$, that is $\theta_{q}(r)=\frac{q^{r+1}-1}{q-1}=q^{r}+q^{r-1}+\ldots+1$ for $r\geq 0$ and $\theta_{q}(-1)=0$. We will use the following result from [17].
Theorem 2.5.
Let $\mathscr{X}$ be a hypersurface of degree $\nu$ in $\mathrm{PG}(r,q)$ without an $\mathbb{F}_{q}$-linear component and let $N_{q}(\mathscr{X})$ be the number of $\mathbb{F}_{q}$-rational points of $\mathscr{X}$. Then
$N_{q}(\mathscr{X})\leq(\nu-1)q^{r-1}+\nu q^{r-2}+\theta_{q}(r-3).$ (4)
3 A family of minimal linear $q$-ary codes, $q$ even
Let $H$ be a quasi-Hermitian variety of $\mathrm{PG}(r,q^{2})$. As shown in [2], the linear code $\mathscr{C}(H)$ generated by $H$ is minimal. This was established by applying Theorem 2.2, since the sufficient condition for minimality given by Ashikhmin and Barg (1) could not be used.
Consider the BT quasi-Hermitian variety $\mathscr{H}_{\varepsilon}^{r}=(\mathscr{V}_{\varepsilon}^{r}\setminus\Pi_{\infty})\cup\mathscr{\mathscr{H}_{\varepsilon,\infty}^{r}}$. We ask whether $\mathscr{V}_{\varepsilon}^{r}$ induces minimal codes.
3.1 The variety $\mathscr{V}_{\varepsilon}^{3}$
We first study the projective linear code $\mathscr{C}^{3}_{\varepsilon}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$. As an initial step, we investigate the variety at infinity, defined as $\mathscr{V}^{3}_{\varepsilon,\infty}:=\mathscr{V}^{3}_{\varepsilon}\cap\Pi_{\infty}$.
Lemma 3.1.
Let $\Pi_{\infty}$ be the plane at infinity of $\mathrm{PG}(3,q^{2})$. Then,
$\mathscr{V}^{3}_{\varepsilon,\infty}=\begin{cases}\ell_{0}&\text{ if }e\equiv 1\;mod\,4\\ \ell_{1}\cup\ell_{2}\cup\ell_{3}&\text{ if }e\equiv 3\;mod\,4,\end{cases}$
where $\ell_{0}:X_{0}=X_{1}+X_{2}=0$ and $\ell_{i}:X_{0}=X_{1}+c_{i}X_{2}=0$ for $i=1,2,3$ and $c_{1}$, $c_{2}$ and $c_{3}$ are the $(2^{\frac{e-1}{2}}+1)$-th roots of unity in $\mathbb{F}_{q^{2}}$.
Proof.
The points in $\mathscr{V}^{3}_{\varepsilon,\infty}$, satisfy equation
$X_{1}^{\sigma+2}+X_{2}^{\sigma+2}=0.$
Recalling that $\sigma:x\longmapsto x^{2^{\frac{e+1}{2}}}$, we obtain
$X_{1}^{2^{\frac{e+1}{2}}+2}+X_{2}^{2^{\frac{e+1}{2}}+2}=X_{1}^{2(2^{\frac{e-1}{2}}+1)}+X_{2}^{2(2^{\frac{e-1}{2}}+1)}=0$
and, hence,
$X_{1}^{2^{\frac{e-1}{2}}+1}+X_{2}^{2^{\frac{e-1}{2}}+1}=0.$
An easy non-trivial solution to this equation is the point $P_{\infty}(0,0,0,1)$. In order to find other solutions, we assume $X_{2}\neq 0$ and study
$\left(\frac{X_{1}}{X_{2}}\right)^{2^{\frac{e-1}{2}}+1}=1.$
This reduces to the study of the $(2^{\frac{e-1}{2}+1})$-th roots of unity in $\mathbb{F}_{q^{2}}$ and, hence, to compute the $\gcd(2^{\frac{e-1}{2}}+1,q^{2}-1)$. Let $d=\gcd\left(\frac{e-1}{2},2e\right)$. Since
$\gcd\left(2^{\frac{e-1}{2}}+1,2^{2e}-1\right)=\begin{cases}1&\text{if $\frac{2e}{d}$ is odd}\\ 2^{d}+1&\text{otherwise}\end{cases}$
and $e$ is odd, then it is enough to compute $d$.
If $e\equiv 1\pmod{4}$, then $e=4m+1$ for some $m\in\mathbb{N}$. So, $d=\gcd(2m,2e)=2\gcd(m,e)$, implying that $2e/d$ is odd. If $e\equiv 3\pmod{4}$, then $e=4m+3$ for some $m\in\mathbb{N}$ and
$d=\gcd(2m+1,2(4m+3))=\gcd(2m+1,4m+3)=1.$
This completes the proof. ∎
By Theorem 2.4, we have that $|\mathscr{V}^{3}_{\varepsilon}\setminus\Pi_{\infty}|=q^{5}$. So, by Lemma 3.1, we get the following.
Theorem 3.2.
[…]
Theorem 3.3.
If $e\equiv 3\pmod{4}$, then the weights of the projective code $\mathscr{C}^{3}_{\varepsilon}:=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ belong to the set
$\large\{q^{5},q^{5}-q^{3}+3q^{2},q^{5}-q^{3}+2q^{2},q^{5}-q^{3}+3q^{2}+q-2,\\ q^{5}-q^{3}+2q^{2}+q-2\large\}.$
Proof.
We have to study the cardinality of the intersections between $\mathscr{V}^{3}_{\varepsilon}$ and a plane $\pi:aX_{0}+bX_{1}+cX_{2}+dX_{3}=0$ of $\mathrm{PG}(3,q^{2})$. If $(b,c,d)=(0,0,0)$ then $\pi=\Pi_{\infty}$ and by Lemma 3.1, $|\mathscr{V}^{3}_{\varepsilon,\infty}|=3q^{2}+1$.
Next, we distinguish the following cases:
  1. a)
    $d=0$, $b\neq 0\neq c$ and $\frac{b}{c}\in\{c_{1},c_{2},c_{3}\}$, where $c_{1},c_{2},c_{3}$ are the $(2^{\frac{e-1}{2}}+1)$-th roots of unity;
  2. b)
    $d=0$, $(b,c)\neq(0,0)$ and $\frac{b}{c}$ is not a $(2^{\frac{e-1}{2}}+1)$-th root of unity;
  3. c)
    $d\neq 0$.
If $\pi$ is of type a) then $|\pi\cap\mathscr{V}^{3}_{\varepsilon,\infty}|=q^{2}+1$ whereas if it is of type b) then $\pi\cap\mathscr{V}^{3}_{\varepsilon,\infty}$ consists only of the point $P_{\infty}(0,0,0,1)$.
In order to determine the number of the affine points of $\mathscr{V}^{3}_{\varepsilon}\cap\pi$ in cases a) or b) we can assume without loss of generality $b=1$. The case of $b=0$ and $c\neq 0$ follows immediately by swapping the roles of $X_{1}$ and $X_{2}$. So we need to compute the number of solutions for the following system in unknowns $(x,y,z)\in\mathbb{F}_{q^{2}}^{3}$
$\begin{cases}z^{q}+z=\Gamma_{\varepsilon}(x)+\Gamma_{\varepsilon}(y)\\ x+cy+a=0.\end{cases}$ (6)
In order to solve the system above, let consider the $\mathbb{F}_{q}$-basis $\{1,\varepsilon\}$ of $\mathbb{F}_{q^{2}}$ introduced in Section 2.2. Then,
$x=x_{0}+\varepsilon x_{1}\;\;y=y_{0}+\varepsilon y_{1}\;\;z=z_{0}+\varepsilon z_{1}\;\;a=a_{0}+\varepsilon a_{1}\;\;c=c_{0}+\varepsilon c_{1}.$
with $x_{i},y_{i},z_{i},a_{i},c_{i}\in\mathbb{F}_{q}$, $i=1,2$. Note that $x+x^{q}=x_{1}$,
$\begin{split}\Gamma_{\varepsilon}(x)&=[(x_{0}+\varepsilon x_{1})+x_{1}\varepsilon]^{\sigma+2}+x_{1}^{\sigma}+x_{1}^{2}\varepsilon+(x_{0}+x_{1}\varepsilon)^{q+1}+x_{0}^{2}+\varepsilon^{2}x_{1}^{2}\\ &=x_{0}^{\sigma+2}+x_{1}^{\sigma}+x_{1}^{2}\varepsilon+x_{0}^{2}+x_{0}x_{1}+\varepsilon^{q+1}x_{1}^{2}+x_{0}^{2}+x_{0}^{2}+\varepsilon^{2}x_{1}^{2}\\ &=x_{0}^{\sigma+2}+x_{1}^{\sigma}+x_{0}x_{1}+x_{1}^{2}(\varepsilon^{2}+\varepsilon+\delta)\\ &=x_{0}^{\sigma+2}+x_{0}x_{1}+x_{1}^{\sigma}.\end{split}$ (7)
and, similarly, $\Gamma_{\varepsilon}(y)=y_{0}^{\sigma+2}+y_{0}y_{1}+y_{1}^{\sigma}.$
Substituting the expressions obtained above in System (6), we obtain
$\begin{cases}z_{1}=x_{0}^{\sigma+2}+x_{0}x_{1}+x_{1}^{\sigma}+y_{0}^{\sigma+2}+y_{0}y_{1}+y_{1}^{\sigma}\\ x_{0}+\varepsilon x_{1}=(c_{0}y_{0}+\delta c_{1}y_{1}+a_{0})+\varepsilon(c_{1}y_{0}+c_{0}y_{1}+c_{1}y_{1}+a_{1})\end{cases}$ (8)
By the second equation in (8), we derive
$x_{0}=c_{0}y_{0}+\delta c_{1}y_{1}+a_{0}\text{ and }x_{1}=c_{1}y_{0}+c_{0}y_{1}+c_{1}y_{1}+a_{1}.$ (9)
We see that the value of $x$ is uniquely determined by $y$. Substituting (9) in the first equation of (6), we see that $z_{1}$ is also uniquely determined by $y$, while $z_{0}$ ranges in $\mathbb{F}_{q}$. Since $y$ ranges in $\mathbb{F}_{q^{2}}$, we get that system (6) has $q^{3}$ solutions; thus the cardinality of $\mathscr{V}^{3}_{\varepsilon}\cap\pi$ is respectively $q^{3}+q^{2}+1$ in case a) or $q^{3}+1$ in case b).
We now study $|\pi\cap\mathscr{V}^{3}_{\varepsilon}|$ in case c). Recall that $|\pi\cap\mathscr{V}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|=|\pi\cap\mathscr{H}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|$.
Since $\mathscr{H}^{3}_{\varepsilon}$ is a quasi Hermitian variety, the cardinality of the intersection between $\mathscr{H}^{3}_{\varepsilon}$ and a generic plane in $\mathrm{PG}(3,q^{2})$ is either $q^{3}+1$ or $q^{3}+q^{2}+1$; so, we have
$\{q^{3}+1,q^{3}+q^{2}+1\}\ni|\pi\cap\mathscr{H}^{3}_{\varepsilon}|=|\pi\cap\mathscr{H}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|+|\pi\cap\mathscr{H}^{3}_{\varepsilon}\cap\Pi_{\infty}|.$
Recalling that the set of infinity points in $\mathscr{H}^{3}_{\varepsilon}$ is given by $q+1$ lines passing through $P_{\infty}(0,0,0,1)$ and $\pi$ does not contain $P_{\infty}$, we get
$|\pi\cap\mathscr{V}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|=|\pi\cap\mathscr{H}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|\in\{q^{3}-q,q^{3}+q^{2}-q\}.$
So, if $\pi$ is of type c), $|\pi\cap\mathscr{V}^{3}_{\varepsilon}|\in\{q^{3}-q+3,q^{3}+q^{2}-q+3\}$. Finally, by (5), the weights of $\mathscr{C}^{3}_{\varepsilon}$ with respect to planes of $\mathrm{PG}(3,q^{2})$ are as in the statement, concluding the proof. ∎
Theorem 3.4.
If $e\equiv 1\pmod{4}$, then the weights of $\mathscr{C}^{3}_{\varepsilon}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ belong to the set
$\large\{q^{5},q^{5}-q^{3},q^{5}-q^{3}+q^{2},q^{5}-q^{3}+q,q^{5}-q^{3}+q^{2}+q\large\}.$
Proof.
Since $\mathscr{V}^{3}_{\varepsilon,\infty}:=\Pi_{\infty}\cap\mathscr{V}^{3}_{\varepsilon}:X_{1}+X_{2}=X_{0}=0$, the weight for the plane at infinity is $|\mathscr{V}^{3}_{\varepsilon}|-q^{2}-1$. We split the set of affine projective planes with equations $aX_{0}+bX_{1}+cX_{2}+dX_{3}=0$, $(b,c,d)\in\mathbb{F}_{q^{2}}^{3}\setminus\{(0,0,0)\}$ of $\mathrm{PG}(3,q^{2})$ in three subsets:
  1. a)
    $d=0$ and $b=c$, then $\pi:aX_{0}+X_{1}+X_{2}=0$ and $|\pi\cap\mathscr{V}^{3}_{\varepsilon,\infty}|=q^{2}+1$;
  2. b)
    $d=0$ and $b\neq c$, then $\pi\cap\mathscr{V}^{3}_{\varepsilon,\infty}=P_{\infty}(0,0,0,1)$;
  3. c)
    $d\neq 0$, then $\pi\cap\mathscr{V}^{3}_{\varepsilon,\infty}=Q$ where $Q\neq P_{\infty}$.
Now we compute the cardinalities of the intersections between $\mathscr{V}^{3}_{\varepsilon}$ and the planes $\pi$ mentioned above. For the plane of type a) the cardinality is $q^{3}+q^{2}+1$ and, consequently the weight is $|\mathscr{V}^{3}_{\varepsilon}|-q^{3}-q^{2}-1$; for the planes of type b) the cardinality is $q^{3}+1$ and the weight is $|\mathscr{V}^{3}_{\varepsilon}|-q^{3}-1$. This is obtained studying the number of solutions of System (6) in the proof of Theorem 3.3. Finally, we tackle the study of $|\pi\cap\mathscr{V}^{3}_{\varepsilon}|$ for a plane of type c). Then, since $\mathscr{V}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})=\mathscr{H}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})$ and $|\pi\cap\mathscr{H}^{3}_{\varepsilon}|\in\{q^{3}+1,q^{3}+q^{2}+1\}$, we have that
$\begin{split}|\pi\cap\mathscr{V}^{3}_{\varepsilon}|&=|\pi\cap\mathscr{V}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|+|\pi\cap\mathscr{V}^{3}_{\varepsilon,\infty}|\\ &=|\pi\cap\mathscr{H}^{3}_{\varepsilon}\cap\mathrm{AG}(3,q^{2})|+1=|\pi\cap\mathscr{H}^{3}_{\varepsilon}|-|\pi\cap\mathscr{H}^{3}_{\varepsilon}\cap\Pi_{\infty}|+1\\ &=|\pi\cap\mathscr{H}^{3}_{\varepsilon}|-q\in\{q^{3}-q+1,q^{3}+q^{2}-q+1\}.\end{split}$ (10)
So, the weight of a plane of type c)) is either $|\mathscr{V}^{3}_{\varepsilon}|-(q^{3}+q^{2}-q+1)$ or $|\mathscr{V}^{3}_{\varepsilon}|-(q^{3}-q+1)$. This concludes the proof. ∎
Theorem 3.5.
The code $\mathscr{C}^{3}_{\varepsilon}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ is a minimal linear code if and only if $e\equiv 3\pmod{4}$.
Proof.
If $e\equiv 3\pmod{4}$, then, by Theorem 3.3, the ratio between the minimum and maximum weights of $\mathscr{C}^{3}_{\varepsilon}$ is
$\frac{|\mathscr{V}^{3}_{\varepsilon}|-q^{3}-q^{2}-1}{|\mathscr{V}^{3}_{\varepsilon}|-3q^{2}-1}=\frac{q^{5}-q^{3}+2q^{2}}{q^{5}}>\frac{q^{2}-1}{q^{2}};$
this proves the minimality by Condition (1). If $e\equiv 1\pmod{4}$, then $\langle\Pi_{\infty}\cap\mathscr{V}^{3}_{\varepsilon}\rangle\neq\Pi_{\infty}$ and hence, from Theorem 2.2, the code $\mathscr{C}^{3}_{\varepsilon}$ is not minimal. ∎
Corollary 3.6.
The code $\mathscr{C}_{\varepsilon}^{3}=\mathscr{C}(\mathscr{V}^{3}_{\varepsilon})$ is a $|\mathscr{V}_{\varepsilon}^{3}|,4]-$linear code with minimum distance $d=|\mathscr{V}_{\varepsilon}^{3}|-q^{3}-q^{2}-1$. Moreover, its $m$-th higher minimum weights $d_{m}$ with $m\in\{1,2,3\}$ are:
$d_{1}=|\mathscr{V}^{3}_{\varepsilon}|-q^{3}-q^{2}-1;\;\;d_{2}=|\mathscr{V}^{3}_{\varepsilon}|-q^{2}-1;\;\;d_{3}=|\mathscr{V}^{3}_{\varepsilon}|-1.$
Remark 3.7.
We stress that in [11], the authors provided a general method to transform a minimal code satisfying the Ashikhmin–Barg condition into a minimal code that violates it.
3.2 The variety $\mathscr{V}_{\varepsilon}^{r}$, $r\geq 4$
In this section we are going to determine the parameters of the linear code $\mathscr{C}^{r}_{\varepsilon}=\mathscr{C}(\mathscr{V}_{\varepsilon}^{r})$ generated by the projective points of $\mathscr{V}^{r}_{\varepsilon}$ in $\mathrm{PG}(r,q^{2})$, $r\geq 4$ showing that it is a minimal code for any odd integer $e$.
Lemma 3.8.
Let $r\geq 2$ and $q=2^{e}$ with $e\geq 3$ odd. Then the Fermat hypersurface $\mathscr{F}_{n}^{r}$ of degree $n=2^{\frac{e-1}{2}}+1$ in $\mathrm{PG}(r,q^{2})$ defined by the equation
$\mathscr{F}^{r}_{n}:X_{0}^{2^{\frac{e-1}{2}}+1}+X_{1}^{2^{\frac{e-1}{2}}+1}+\ldots+X_{r}^{2^{\frac{e-1}{2}}+1}=0$
spans $\mathrm{PG}(r,q^{2})$, that is $\langle\mathscr{F}^{r}_{n}\rangle=\mathrm{PG}(r,q^{2})$. When $e=1$ the hypersurface $\mathscr{F}_{n}^{r}$ is a hyperplane in $\mathrm{PG}(r,4)$.
Proof.
If $e=1$, then $\mathscr{F}^{r}_{n}$ is a degenerate quadric in characteristic $2$, splitting in a hyperplane counted twice. Assume now $e\geq 2$ and set $n=2^{\frac{e-1}{2}}+1$. The proof is by induction on $r$. Suppose $r=2$. Since $\mathscr{F}^{2}_{n}$ is non-singular and absolutely irreducible, the Hasse-Weil bound assures that the number $N_{q^{2}}:=N_{q^{2}}(\mathscr{F}^{r}_{n})$ of its $\mathbb{F}_{q^{2}}$-rational points is at least $q^{2}+1-(n-1)(n-2)q$. Since each line of the projective plane meets $\mathscr{F}^{2}_{n}$ in at most $n$ points and $N_{q^{2}}>n$ then the rational points on $\mathscr{F}^{2}_{n}$ generate the plane. Suppose now that the points of $\mathscr{F}^{r-1}_{n}$ generate $\mathrm{PG}(r-1,q^{2})$ and consider $\mathscr{F}^{r}_{n}\cap\Pi$, where $\Pi$ is the hyperplane with equation $X_{r}=0$ of $\mathrm{PG}(r,q^{2})$. Then $\mathscr{F}^{r}_{n}\cap\Pi=\mathscr{F}^{r-1}_{n}$; so, by induction, $\mathscr{F}^{r-1}_{n}$ generates the hyperplane $\Pi$. On the other hand, the point $M(1,0,\ldots,0,1)$ belongs to $\mathscr{F}^{r}_{n}\setminus\Pi$ and hence
$\mathrm{PG}(r,q^{2})=\langle M,\Pi\rangle=\langle M,\mathscr{F}^{r-1}_{n}\rangle\subseteq\langle\mathscr{F}_{n}^{r}\rangle.$
Theorem 3.9.
For $r\geq 4$ and $e\geq 3$ odd, the set of rational points $\mathscr{V}^{r}_{\varepsilon}\subset\mathrm{PG}(r,q^{2})$ is cutting, i.e. for any hyperplane of $\mathrm{PG}(r,q^{2})$,
$\langle\Pi\cap\mathscr{V}^{r}_{\varepsilon}\rangle=\Pi.$ (11)
Proof.
Firstly, we prove that $\langle\mathscr{V}^{r}_{\varepsilon}\cap\Pi_{\infty}\rangle=\Pi_{\infty}$. Put $\mathscr{V}^{r}_{\varepsilon,\infty}=\mathscr{V}^{r}_{\varepsilon}\cap\Pi_{\infty}$. By Equation (2), $\mathscr{V}^{r}_{\varepsilon,\infty}$ has equations
$X_{0}=X_{1}^{2^{\frac{e-1}{2}}+1}+X_{2}^{2^{\frac{e-1}{2}}+1}+\ldots+X_{r-1}^{2^{\frac{e-1}{2}}+1}=0.$
Then, this variety is a cone through the point $P_{\infty}(0,0,\ldots,1)$ and a Fermat hypersurface $\mathscr{F}^{r-2}_{n}$, $n=2^{\frac{e-1}{2}}+1$, contained in the (r2)(r-2)-dimensional subspace $\Pi_{r-2}:X_{0}=X_{r}=0$. By Lemma 3.8, this variety spans $\Pi_{r-2}$ for $e\geq 3$. So, since $P_{\infty}$ belongs to $\mathscr{V}^{r}_{\varepsilon,\infty}\setminus\Pi_{r-2}$, we have
$\langle\mathscr{V}^{r}_{\varepsilon,\infty}\rangle=\langle P_{\infty},\mathscr{V}^{r}_{\varepsilon,\infty}\cap\Pi_{r-2}\rangle=\langle P_{\infty},\Pi_{r-2}\rangle=\Pi_{\infty}.$
In order to prove (11), for any other hyperplane $\Pi$ distinct from $\Pi_{\infty}$, it is enough to note that
$|\mathscr{V}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|=|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|,$
and to show that the size of the set $\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi$ is strictly greater than $|\mathrm{PG}(r-2,q^{2})|=\theta_{q^{2}}(r-2)$ points.
We will use the following property, see [1]. Let $\mathscr{H}^{r}:=\mathscr{H}(r,q^{2})$ be the Hermitian variety of equation $\mathscr{H}^{r}:X_{r}^{q}+X_{r}=X_{1}^{q+1}+\ldots+X_{r-1}^{q}$ and $\Pi$ be a hyperplane through the point $P_{\infty}$. Then
$\mathscr{H}^{r}\cap\Pi\cap\Pi_{\infty}=\mathscr{H}_{\varepsilon}^{r}\cap\Pi\cap\Pi_{\infty}.$ (12)
We split the discussion in four cases:
  1. i)
    $P_{\infty}\not\in\Pi$ and $|\mathscr{H}^{r}_{\varepsilon}\cap\Pi|=|\mathscr{H}(r-1,q^{2})|$ see [1]. In this case $\mathscr{H}^{r}_{\varepsilon}\cap\Pi_{\infty}\cap\Pi$ is an $\mathscr{H}(r-2,q^{2})$, otherwise $P_{\infty}$ would belong to $\Pi$ against the hypothesis, cf. (3). Hence,
    $\begin{split}|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|&=|\mathscr{H}(r-1,q^{2})|-|\mathscr{H}(r-2,q^{2})|\\ &=q^{r-2}(q^{r-1}-(-1)^{r-1}).\end{split}$ (13)
    Let us consider
[…]
    In order to get the claim, it is enough to show that $q^{r}-q^{r-2}-q^{r-1}-1$ is positive for any $q\geq 2$ and $r\geq 4$.
    Since $q^{r-\ell}\leq\frac{q^{r}}{2^{\ell}}$ for any $q\geq 2$, $r\geq 4$ and $\ell\in\{0,\ldots,r\}$ and by (14), we have
    $q^{r}-q^{r-2}-q^{r-1}-1\geq\frac{q^{r}}{4}-1>0.$
  2. ii)
    $P_{\infty}\not\in\Pi$ and $|\Pi\cap\mathscr{H}^{r}_{\varepsilon}|=|\Pi_{0}\mathscr{H}(r-2,q^{2})|$. Since $P_{\infty}\notin\Pi$, we have that $\mathscr{H}^{r}_{\varepsilon}\cap\Pi_{\infty}\cap\Pi$ is again an $\mathscr{H}(r-2,q^{2})$. We proceed as in point i):
    $\begin{split}|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|&=|\Pi_{0}\mathscr{H}(r-2,q^{2})|-|\mathscr{H}(r-2,q^{2})|=\\ &=(-1)^{r-1}q^{r-1}+\frac{(q^{r-1}-(-1)^{r-1})(q^{r}-q^{r-2})}{(q^{2}-1)}\\ &=(-1)^{r-1}q^{r-1}+q^{r-2}(q^{r-1}-(-1)^{r-1}).\end{split}$ (15)
    Studying the case for $r$ even, we obtain
    $|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|=q^{2r-3}-q^{r-1}+q^{r-2}=q^{r-2}(q^{r-1}-q+1)$
    and hence,
    $(q^{2}-1)\left(|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|-\theta_{q^{2}}(r-2)\right)=\\ (q^{2}-1)q^{r-2}(q^{r-1}-q+1)-(q^{2(r-1)}-1).$
    Then, we have that
    $\begin{split}q^{2r-1}&+q^{r}-q^{r+1}-q^{2r-3}-q^{r-2}+q^{r-1}-q^{2(r-1)}+1\\ &>q^{r-2}(q^{r+1}+q^{2}-q^{3}-q^{r-1}-1+q-q^{r})\\ &>q^{r-2}(q^{r+1}+q^{2}-q^{3}-q^{r-1}-q^{r})\\ &>q^{r-2}(q^{r-1}(q^{2}-q-1)+q^{2}-q^{3})\\ &\geq q^{r}(q^{r-3}+1-q)>0,\end{split}$ (16)
    for any $r\geq 4$ and $q\geq 2$. Then,
    $|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|>\theta_{q^{2}}(r-2).$
    Similarly in the case of $r$ odd we have that
    $|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|=q^{r-1}+q^{2r-3}-q^{r-2}$
    and therefore
    $\begin{split}(q^{2}-1)&\left(|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|-\theta_{q^{2}}(r-2)\right)\\ &>q^{r-2}(q^{r+1}-q^{r}-q^{r-1}+q^{3}-q^{2}-q+1)\\ &>q^{r-2}(q^{r-1}(q^{2}-q-1)+q(q^{2}-q-1)+1)\\ &>q^{r-1}(q^{r-2}+1)(q^{2}-q-1)>0\end{split}$ (17)
    for any $q\geq 2$.
  3. iii)
    $P_{\infty}\in\Pi$ and $|\Pi\cap\mathscr{H}^{r}_{\varepsilon}|=|\Pi_{0}\mathscr{H}(r-2,q^{2})|$.
    We know that $|\mathscr{H}^{r}\cap\Pi|=|\mathscr{H}_{\varepsilon}^{r}\cap\Pi|$ and in particular the hyperplane $\Pi$ is tangent to $\mathscr{H}^{r}$ at some point $Q$ different from $P_{\infty}$. By (12), $\mathscr{H}^{r}\cap\Pi\cap\Pi_{\infty}=\mathscr{H}_{\varepsilon}^{r}\cap\Pi\cap\Pi_{\infty}$ hence, it follows that $\mathscr{H}_{\varepsilon}^{r}\cap\Pi\cap\Pi_{\infty}=\mathscr{H}^{r}\cap P_{\infty}^{\perp}\cap Q^{\perp}$ with $P_{\infty}\in Q^{\perp}$; so $|\mathscr{H}_{\varepsilon}^{r}\cap\Pi_{\infty}\cap\Pi|=|\Pi_{1}\mathscr{H}(r-4,q^{2})|$.
    Therefore, in order to prove that $\langle\mathscr{V}^{r}_{\varepsilon}\cap\Pi\rangle=\Pi$, it is enough to show
    $|\Pi_{0}\mathscr{H}(r-2,q^{2})|-|\Pi_{1}\mathscr{H}(r-4,q^{2})|>\theta_{q^{2}}(r-2).$
    In case of $r$ odd we have
    We observe that
    $|\Pi_{0}\mathscr{H}(r-2,q^{2})|-|\Pi_{1}\mathscr{H}(r-4,q^{2})|=\\ \frac{(q^{r}+(-1)^{r-1})(q^{r-1}-(-1)^{r-1}}{q^{2}-1}+\\ (-1)^{r-1}q^{r-1}-\frac{(q^{r-3}-(-1)^{r-4})(q^{r-4}-(-1)^{r-4})q^{4}}{q^{2}-1}-q^{2}-1$
    which is equal to $q^{2r-3}$ for $r$ even as well as for $r$ odd.
    Now,
    $(q^{2}-1)(|\Pi_{0}\mathscr{H}(r-2,q^{2})|-|\Pi_{1}\mathscr{H}(r-4,q^{2})|-\theta_{q^{2}}(r-2))\\ =q^{2r-1}-q^{2r-3}-q^{2r-2}+1=q^{2r-3}(q^{2}-q-1)+1$
    where the last term is clearly positive for $q\geq 2$.
  4. iv)
    $P_{\infty}\in\Pi$, $|\Pi\cap\mathscr{H}^{r}_{\varepsilon}|=|\mathscr{H}(r-1,q^{2})|$.
    In this case $|\Pi\cap\Pi_{\infty}\cap\mathscr{H}^{r}_{\varepsilon}|=|\Pi_{0}\mathscr{H}(r-3,q^{2})|$. Then,
    $|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|=\\ |\mathscr{H}(r-1,q^{2})|-|\Pi_{0}\mathscr{H}(r-3,q^{2})|=\frac{(q^{r}+(-1)^{r-1})(q^{r-1}-(-1)^{r-1})}{q^{2}-1}\\ -\frac{(q^{r-1}+(-1)^{r-2})(q^{r-2}-(-1)^{r-2})}{q^{2}-1}-(-1)^{r-2}q^{r-2}$ (18)
    Both for $r$ even and odd, we get
    $|\mathscr{H}^{r}_{\varepsilon}\cap\mathrm{AG}(r,q^{2})\cap\Pi|=|\mathscr{H}(r-1,q^{2})|-|\Pi_{0}\mathscr{H}(r-3,q^{2})|=q^{2r-3}.$
    Thus, we can argue as in the case above.
Theorem 3.10.
Let $\mathscr{V}^{r}_{\varepsilon}\subset\mathrm{PG}(r,q^{2})$, $r\geq 4$. Then,
$|\mathscr{V}^{r}_{\varepsilon}|=q^{2r-1}+|\mathscr{F}^{r-2}_{n}|q^{2}+1,$
where $n=2^{\frac{e-1}{2}}+1$.
Proof.
The result follows observing that $\mathscr{V}^{r}_{\varepsilon}\cap\Pi_{\infty}$ is a cone with vertex the point $P_{\infty}(0,0,\ldots,1)$ over a Fermat hypersurface of the (r2)(r-2)-dimensional subspace with equations $X_{0}=X_{r}=0$. ∎
By Theorem 3.10, in order to get the length and the weights of the projective linear code $\mathscr{C}_{\varepsilon}^{r}$, $r\geq 4$, it is necessary to determine the size of the Fermat hypersurface $\mathscr{F}_{n}^{r-2}$, $r\geq 4$ and $n=2^{\frac{e-1}{2}}+1$, lying in the (r2)(r-2)-dimensional subspace of $\mathrm{PG}(r,q^{2})$ of equations $X_{0}=X_{r}=0$. Thus, we prove the following.
Proposition 3.11.
The number $N_{q^{2}}$ of $\mathbb{F}_{q^{2}}$-rational points of $\mathscr{F}^{r}_{n}$ in $\mathrm{PG}(r,q^{2})$, $r\geq 2$, satisfies the following properties:
  1. (i)
    if $e\equiv 1\pmod{4}$ then $N_{q^{2}}=\theta_{q^{2}}(r-1)$;
  2. (ii)
    if $e\equiv 3\pmod{4}$ then
    $N_{q^{2}}\leq(n-1)q^{2(r-1)}+nq^{2(r-2)}+\theta_{q^{2}}(r-3).$
  3. (iii)
    if $e=3$ and $r=2$ then $N_{q^{2}}=(q+1)^{2}.$
Proof.
  1. (i)
    If $e=1$, then $\mathscr{F}_{n}^{r}$ is a hyperplane in $\mathrm{PG}(r,q^{2})$ and the statement holds. Suppose now $e\geq 3$. The proof is by induction on $r$. For $r=2$, $\mathscr{F}^{2}_{n}$ is the plane curve of equation $X_{0}^{n}+X_{1}^{n}+X_{2}^{n}=0$. Then, the points of $\mathscr{F}^{2}_{n}$ are of the following forms:
    1. (a)
      $(0,1,x_{2})$, $x_{2}^{n}=1$;
    2. (b)
      $(1,x_{1},0)$, $x_{1}^{n}=1$;
    3. (c)
      $(1,x_{1},x_{2})$, $x_{1}^{n}=\alpha\neq 1$ and $x_{2}^{n}=1+\alpha\neq 0$, where $\alpha\in\mathbb{F}_{q^{2}}$.
    Since $\gcd(q^{2}-1,n)=\gcd(2^{2e}-1,2^{\frac{e-1}{2}}+1)=1$ for $e\equiv 1\pmod{4}$, then any $m\in\mathbb{F}_{q^{2}}\setminus\{0\}$ is an $n$-th power of an element in $\mathbb{F}_{q^{2}}$. Furthermore, there is precisely one point of type (i)(a) and one point of type (i)(b). As for type (i)(c), $x_{1}^{n}=1$ implies $x_{1}=1$; so there are exactly $q^{2}-1$ possible second coordinates for any point of type (i)(c). Given any such second coordinate, the third coordinate $x_{2}$ is uniquely determined. So, there are precisely $q^{2}-1$ points of type (i)(c). So, $\mathscr{F}^{2}_{n}$ has $q^{2}+1$ rational points. Now, by the induction hypothesis, let suppose that the points of $\mathscr{F}^{r-1}_{n}$ are $\theta_{q^{2}}(r-2)=q^{2(r-2)}+q^{2(r-3)}+\ldots+q^{2}+1$ in total and consider the hypersurface $\mathscr{F}^{r}_{n}:\ X_{0}^{n}+X_{2}^{n}+\ldots+X_{r}^{n}=0$ of $\mathrm{PG}(r,q^{2})$. The points of $\mathscr{F}^{r}_{n}$ for which $x_{r}=0$ are $\theta_{q^{2}}(r-2)$ by induction, whereas the points of our hypersurface for which $x_{r}=1$ are $q^{2(r-1)}$ in total; so the result follows.
  2. (ii)
    This is a consequence of Theorem 2.5 applied to the hypersurface $\mathscr{F}^{r}_{n}$.
  3. (iii)
    In this case, the plane curve $\mathscr{F}^{2}_{3}$ turns out to be maximal by Theorem [15, Theorem 10.65]. So the result follows from the Hasse-Weil bound.
Theorem 3.12.
The linear code $\mathscr{C}^{r}_{\varepsilon}$ generated by the projective points of $\mathscr{V}^{r}_{\varepsilon}$ in $\mathrm{PG}(r,q^{2})$, $r\geq 4$ is a $(r+1)$-dimensional minimal code for any odd integer $e$ with length $|\mathscr{V}_{\varepsilon}^{r}|$.
Proof.
The thesis follows from Theorem 2.2 and Theorem 3.9. ∎
Now, consider the multiset $\overline{\mathscr{V}_{\varepsilon}^{r}}$ in $\mathrm{PG}(r,q^{2})$, $r\geq 3$, that consists of the points in $(\mathscr{V}_{\varepsilon}^{r}\setminus\Pi_{\infty})$ plus the point $P_{\infty}$, which is assigned multiplicity $j\geq 1$. Using an approach similar to that used in [3] we get the following result.
Theorem 3.13.
The linear code associated with $\overline{\mathscr{V}_{\varepsilon}^{r}}$ is a $q^{2r-1}+j,r+1]_{q^{2}}$ code with weights:
$q^{2r-1},q^{2r-1}-q^{2r-3},q^{2r-1}-q^{2r-3}+(-1)^{r-1}q^{r-2}+j,$
$q^{2r-1}-q^{2r-3}+(-1)^{r-1}q^{r-2}-(-1)^{r-1}q^{r-1}+j.$
Furthermore, it is a $3$-weight code in the following cases
  1. (i)
    $r$ odd and $j=q^{r-1}-q^{r-2}$ or $j=q^{2r-3}-q^{r-2}$ or $j=q^{2r-3}+q^{r-1}-q^{r-2}$;
  2. (ii)
    $r$ even and $j=q^{r-2}$, or $j=q^{2r-3}+q^{r-2}$ or $j=q^{2r-3}-q^{r-1}+q^{r-2}$.
Proof.
It follows from the proofs of Theorem 3.3, Theorem 3.4 and Theorem 3.9. ∎
4 The cutting gap
Minimal projective linear codes and cutting blocking sets with respect to the hyperplanes are essentially the same kind of object, see [7] and [22]. In [13], the notion of $t$-fold strong blocking set $\Lambda$ is introduced, as a set of points in a projective space $\mathrm{PG}(n,q)$ such that every (t1)(t-1)-dimensional projective subspace is spanned by $t$ points in $\Lambda$. Clearly, being a cutting blocking set with respect to the hyperplanes and being a $n$-fold strong blocking set mean the same.
In this section we introduce the concept of cutting gap, as a measure of the amount in which a set $\Lambda$ fails to be a $t$-fold strong blocking set for some $t$. It provides a quantitative measure of how effectively a point set intersects $k$-dimensional subspaces in a projective space. This notion is closely related to the geometry of higher Hamming weights and has applications in coding theory and incidence geometry, as first explored in [13]. In particular, cutting gaps allow us to distinguish between point sets that merely intersect subspaces and those that span them, offering a refined tool for analyzing the structure of algebraic varieties.
Throughout this section, we use projective dimensions and adopt the convention $\dim(\emptyset)=-1$.
Definition 4.1.
Let $\Omega$ be a non-empty set of points in $\mathrm{PG}(r,\mathbb{K})$, and let $0\leq k\leq r$. The $k$-th cutting gap of $\Omega$ is defined as
$\tau_{k}(\Omega)=k-\min\left\{\dim\left(\langle\Pi\cap\Omega\rangle\right)\colon\dim(\Pi)=k\right\}.$
According to [13], $\Omega$ is a $(k+1)$-fold strong blocking set if and only if $\tau_{k}(\Omega)=0$. In this paper, we refer to this property as being $k$-cutting, and we say that $\Omega$ is a $k$-cutting set if $\tau_{k}(\Omega)=0$. For explicit constructions of $k$-cutting sets for selected values of $k$, see [13, §3].
Note that $\Omega$ is a classical cutting set with respect to hyperplanes precisely when $\tau_{r-1}(\Omega)=0$. More generally, $\Omega$ is $k$-cutting if every $k$-dimensional subspace is spanned by its intersection with $\Omega$.
Theorem 4.2.
Let $\Omega$ be a $k$-cutting set in $\mathrm{PG}(r,\mathbb{K})$. Then $\Omega$ is also $\ell$-cutting for all $k\leq\ell\leq r$.
Proof.
If $k=r$, the statement is trivial. Otherwise, suppose $\tau_{k}(\Omega)=0$. Any $(k+1)$-dimensional subspace $\Pi$ can be expressed as the union of its $k$-dimensional subspaces:
$\Pi=\bigcup_{\begin{subarray}{c}\Sigma\leq\Pi\\ \dim(\Sigma)=k\end{subarray}}\Sigma=\bigcup_{\begin{subarray}{c}\Sigma\leq\Pi\\ \dim(\Sigma)=k\end{subarray}}\langle\Sigma\cap\Omega\rangle\subseteq\langle\Pi\cap\Omega\rangle=\Pi.$
Hence, $\tau_{k+1}(\Omega)=0$, and the result follows by induction. ∎
We now compute the cutting gaps for non-degenerate Hermitian varieties using properties of the induced polarity.
Theorem 4.3.
Let $\mathscr{H}^{r}=\mathscr{H}(r,q^{2})$ be a non-degenerate Hermitian variety in $\mathrm{PG}(r,q^{2})$. Then,
$\tau_{r-t}(\mathscr{H}^{r})=\begin{cases}0&\text{if }t>\left\lfloor\frac{r}{2}\right\rfloor,\\ 1&\text{if }t\leq\left\lfloor\frac{r}{2}\right\rfloor.\end{cases}$
Proof.
Hermitian varieties over finite fields always contain at least one rational point. Thus, any subspace $\Pi=\mathrm{PG}(V_{k+1})$ of dimension $k$ in $\mathrm{PG}(r,q^{2})$ intersects $\mathscr{H}^{r}$ in a blocking set for the lines of $\Pi$, which either spans $\Pi$ or is a hyperplane of it.
Let $H_{r+1}$ be the non-singular Hermitian matrix defining $\mathscr{H}^{r}$. For any subspace $\Pi$ of codimension $t$, the induced Hermitian form has rank at least r+12tr+1-2t, as it corresponds to a $(r+1-t)\times(r+1-t)$ minor of $H_{r+1}$. If r+12t>1r+1-2t>1, i.e., $t<r/2$, the form has rank at least $2$, and $\Pi\cap\mathscr{H}^{r}$ spans $\Pi$, so $\tau_{r-t}(\mathscr{H}^{r})=0$.
Conversely, if $t\geq r/2$, there exist subspaces $\Pi$ such that $\Pi\cap\mathscr{H}^{r}$ is a hyperplane of $\Pi$, implying $\tau_{r-t}(\mathscr{H}^{r})=1$. ∎
We leave as an open problem whether the list of cutting gaps in Theorem 4.3 characterizes Hermitian varieties among quasi-Hermitian ones.
In $\mathrm{PG}(2,q)$, the only relevant value is $k=1$. Then: $\tau_{1}(\Omega)=0$ if and only if $\Omega$ is a $2$-fold blocking set for the lines; $\tau_{1}(\Omega)=1$ if $\Omega$ is a blocking set but not $2$-fold; $\tau_{1}(\Omega)=2$ if there exist lines disjoint from $\Omega$. This provides lower bounds on the size of $\Omega$.
A trivial observation is that $\tau_{k}(\Omega)>k$ if and only if there exist $k$-subspaces disjoint from $\Omega$. In such cases, $\tau_{k}(\Omega)$ does not offer any further geometric insight. To refine this, we introduce the notion of modified cutting gaps.
Let $1\leq s<k$. Define:
$\tau_{k,s}(\Omega)=k-\min\left\{\dim\left(\langle\Pi\cap\Omega\rangle\right)\colon\dim(\Pi)=k,\dim(\langle\Pi\cap\Omega\rangle)\geq s\right\}.$
By definition, $\tau_{k,-1}(\Omega)=\tau_{k}(\Omega)$. Moreover, $\tau_{k,s}(\Omega)\leq\tau_{k,s^{\prime}}(\Omega)$ whenever $s>s^{\prime}$, so the sequence is non-increasing in $s$ for fixed $k$. If $\Omega$ spans at least an $s$-space, then $\tau_{k,s}(\Omega)\leq k-s$. While equality is expected in general, the following example shows that this is not always the case.
Example 4.4.
Let $\Omega:=\mathscr{Q}^{-}(3,q)\setminus\{P\}$ be the set of points of an elliptic quadric in $\mathrm{PG}(3,q)$ with one point $P$ removed. Then: $\tau_{2}(\Omega)=\tau_{2,-1}(\Omega)=3$, since the tangent plane at $P$ is disjoint from $\Omega$; $\tau_{2,0}(\Omega)=2$, as tangent planes at other points intersect $\Omega$ in a single point; $\tau_{2,1}(\Omega)=0$, since any plane intersecting $\Omega$ in at least two points meets it in a conic or a conic minus $P$.
We leave to future work a deeper investigation of the properties of modified cutting gaps and the classification of point sets for which $\tau_{k,s}(\Omega)<k-s$ for some values of $k$ and $s$.
4.1 Intersection of $\mathscr{V}^{4}_{\varepsilon}$ with lines
We now study the intersections of $\mathscr{V}^{4}_{\varepsilon}$ with lines, as to show that $\tau_{1}(\mathscr{V}^{4}_{\varepsilon})=2$, i.e. $\mathscr{V}^{4}_{\varepsilon}$ is not a blocking set for the lines of $\mathrm{PG}(4,q^{2})$.
Let $\Pi_{0}$ be the plane of equation $X_{0}=X_{4}=0$. As seen before, $\Pi_{0}\cap\mathscr{V}^{4}_{\varepsilon}$ is the Fermat curve of $\Pi_{0}$ with equation:
$\mathscr{F}^{2}_{s+1}:X_{1}^{s+1}+X_{2}^{s+1}+X_{3}^{s+1}=0.$ (19)
where $s=2^{\frac{e-1}{2}}$. In the following, we are interested in determining an external line to $\mathscr{V}^{4}_{\varepsilon}$; so, it is enough to find a line in $\Pi_{0}$ that does not intersect $\mathscr{F}^{2}_{s+1}$. We split the discussion depending on the value of integer $e$ modulo $4$.
4.1.1 Case $e\equiv 1\pmod{4}$, $e\neq 1$
Let $\ell$ be a line in the plane $\Pi_{0}$ with equations of the form $X_{0}=\alpha X_{1}+\beta X_{2}+X_{3}=X_{4}=0$. Our aim is to prove that there is at least a pair $(\alpha,\beta)\in\mathbb{F}_{q^{2}}^{2}\setminus\{(0,0)\}$ such that $\ell\cap\mathscr{F}^{2}_{s+1}=\emptyset$ or, equivalently that the following system in $X_{1},X_{2},X_{3}$ has no non-trivial solutions:
$\begin{cases}X_{1}^{s+1}+X_{2}^{s+1}+X_{3}^{s+1}=0\\ \alpha X_{1}+\beta X_{2}+X_{3}=0.\end{cases}$ (20)
Substituting the second equation of System (20) in the first, we obtain
$(1+\alpha^{s+1})X_{1}^{s+1}+\alpha^{s}\beta X_{1}^{s}X_{2}+\alpha\beta^{s}X_{1}X_{2}^{s}+(1+\beta^{s+1})X_{2}^{s+1}=0.$
Proposition 4.5.
Let $e\equiv 1\pmod{4}$, $e>1$ and $s=2^{\frac{e-1}{2}}$. Then for every $\alpha\notin\mathbb{F}_{4}$ there exists $1\neq\beta\in\mathbb{F}_{q^{2}}$ such that the equation
$(1+\alpha^{s+1})X_{1}^{s+1}+\alpha^{s}\beta X_{1}^{s}X_{2}+\alpha\beta^{s}X_{1}X_{2}^{s}+(1+\beta^{s+1})X_{2}^{s+1}=0.$ (21)
has no non-trivial solutions in $\mathbb{F}_{q^{2}}^{2}$.
Proof.
Since $X_{2}=0$ gives the trivial solution, we can assume $X_{2}\neq 0$. Moreover, since $\alpha\neq 1$, substituting $X=X_{1}/X_{2}$ in (21), we get
$X^{s+1}+\frac{\alpha^{s}\beta}{1+\alpha^{s+1}}X^{s}+\frac{\alpha\beta^{s}}{1+\alpha^{s+1}}X+\frac{1+\beta^{s+1}}{1+\alpha^{s+1}}=0.$
Putting
$t=\frac{\alpha^{s}\beta}{1+\alpha^{s+1}},\qquad u=\frac{\alpha\beta^{s}}{1+\alpha^{s+1}},\,\,\textnormal{ and }\,\,v=\frac{1+\beta^{s+1}}{1+\alpha^{s+1}},$ (22)
the equation
$X^{s+1}+tX^{s}+uX+v=0$ (23)
has no solution in $\mathbb{F}_{q^{2}}$ if and only if Equation (21) has no non-trivial solution in $\mathbb{F}_{q^{2}}^{2}$. Consider the polynomial $f_{a}=X^{s+1}+X+a$ with $a\in\mathbb{F}_{q^{2}}$. By [6, Theorem 5.6], the set $N_{0}=\{a\in\mathbb{F}_{q^{2}}\colon f_{a}\textnormal{ has no roots in }\mathbb{F}_{q^{2}}\}$ has size $\frac{2}{5}(2^{2e}+1)\geq 410$ for $e\geq 5$. Then, given $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{4}$, we may choose $a\in N_{0}$ such that
$a\neq\frac{(\alpha^{s+1}+1)^{\frac{(s+1)^{2}}{s}-2}}{(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}}\,\,\textnormal{ and }\,\,a\neq\frac{\alpha^{s+1}(\alpha^{s+1}+1)^{\frac{(s+1)^{2}}{s}-2}}{(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}}.$ (24)
Take now $\beta\in\mathbb{F}_{q^{2}}$ such that
$\beta^{s+1}=\frac{(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-1}}{a(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-2}}.$ (25)
Since every element of $\mathbb{F}_{q^{2}}$ is always an $(s+1)$-th power, such an element $\beta$ exists. Firstly, it is straightforward to see that $\beta$ cannot be equal to $1$, otherwise $a$ is equal to the second expression in (24), getting a contradiction. Then, we will show that for this choice of $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{4}$ and $\beta$ as in (25), Equation (23) can be transformed in $X^{s+1}+X+a=0$ and hence it does not have any solutions in $\mathbb{F}_{q^{2}}$. Then, we prove that $u\neq t^{s}$ and $v\neq ut$. Indeed, if $u=t^{s}$, we get
$\frac{\alpha\beta^{s}}{1+\alpha^{s+1}}=\frac{\alpha^{s^{2}}\beta^{s}}{1+\alpha^{s^{2}+s}}.$ (26)
This is equivalent to $\alpha^{s^{2}}=\alpha$ and hence $\alpha^{2^{e-1}-1}=1$. Since $\gcd(e-1,2e)=2$, $\alpha\in\mathbb{F}_{4}$ against the hypotheses. On the other hand, if $v=ut$, we have
$\frac{\beta^{s+1}+1}{\alpha^{s+1}+1}=\frac{\alpha^{s+1}\beta^{s+1}}{(\alpha^{s+1}+1)^{2}}$ (27)
and hence $\alpha^{s+1}+\beta^{s+1}+1=0$. Substituting $\beta^{s+1}=\alpha^{s+1}+1$ in (25), we get
$a(\alpha+\alpha^{s^{2}})^{\frac{(s+1)^{2}}{s}}(\alpha^{s+1}+1)=0.$
Since $\alpha\not\in\mathbb{F}_{4}$, this implies that $a=0$, a contradiction because $a\in N_{0}$. By performing in (21) the substitution $X\longmapsto(u+t^{s})^{s^{-1}}X+t$, we get
$X^{s+1}+X+\frac{tu+v}{(u+t^{s})^{\frac{s+1}{s}}}=0.$ (28)
In order to complete the proof, it is enough to prove that $\frac{tu+v}{(u+t^{s})^{\frac{s+1}{s}}}=a$. Then,
$\frac{tu+v}{(u+t^{s})^{\frac{s+1}{s}}}=\frac{\alpha^{s+1}+\beta^{s+1}+1}{(\alpha^{s+1}+1)^{2}}\cdot\frac{(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}}}{\beta^{s+1}(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}}.$
Substituting (25) in the expression above, we get
$\begin{split}&\frac{(\alpha^{s+1}+1)\left(a(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-2}\right)+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-1}}{(\alpha^{s+1}+1)^{2}\left(a(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-2}\right)}\cdot\\ &\frac{\left(a(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-2})\right)(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}}}{(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-1}}\\ &=\frac{(\alpha^{s+1}+1)^{2}\left(a(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}-2}\right)+(1+\alpha^{s+1})^{\frac{(s+1)^{2}}{s}}}{(\alpha+\alpha^{s^{2}})^{\frac{s+1}{s}}(1+\alpha^{s+1})^{2}}=a.\end{split}$
4.1.2 Case $e\equiv 3\pmod{4}$
Suppose now $e\equiv 3\pmod{4}$ and recall that $\gcd(q^{2}-1,s+1)=3$. We show that in this case there is at least one line of equation $\ell_{\alpha}:X_{0}=\alpha X_{1}+X_{2}=X_{4}=0$ that does not meet the Fermat curve $\mathscr{F}^{2}_{s+1}$. Indeed, the points in $\ell_{\alpha}\cap\mathscr{F}^{2}_{s+1}$ satisfy the system
$\begin{cases}X_{0}=0\\ \alpha X_{1}+X_{2}=0\\ X_{4}=0\\ X_{1}^{s+1}+X_{2}^{s+1}+X_{3}^{s+1}=0.\end{cases}$ (29)
Hence we need to study the solutions of equation
$(\alpha^{s+1}+1)X_{1}^{s+1}+X_{3}^{s+1}=0.$
If $\alpha^{s+1}=1$, then $\alpha\in\{c_{0}=1,c_{1},c_{2}\}$ with $c_{i}^{s+1}=1$, $i=0,1,2$. By the equation above, $X_{3}=0$ and the line $\ell_{c_{i}}$ meets the curve only at the point $(0,1,c_{i},0,0)$. If $\alpha^{s+1}\neq 1$, we can suppose $X_{3}\neq 0$ and consequently also $X_{1}\neq 0$. Putting $X=X_{3}/X_{1}$, we study the solutions of the equation
$X^{s+1}={\alpha^{s+1}+1}$ (30)
for $\alpha$ ranging in $\mathbb{F}_{q^{2}}\setminus\{1,c_{1},c_{2}\}$.
Assume by contradiction that Equation (30) has at least one solution for every $\alpha\in\mathbb{F}_{q^{2}}\setminus\{1,c_{1},c_{2}\}$. If follows that if $\gamma$ is a solution of (30) for a given $\alpha$, then $c_{i}\gamma$ is a solution as well and the corresponding line $\ell_{\alpha}$ has at least three points in common with $\mathscr{F}^{2}_{s+1}$. Moreover, the line with equations $X_{0}=X_{2}=X_{4}=0$ meets the Fermat curve in the points $(0,1,0,1,0)$, $(0,c_{1},0,1,0)$, $(0,c_{2},0,1,0)$. This implies that the number $N$ of rational points of $\mathscr{F}^{2}_{s+1}$ is $N\geq 3(q^{2}-2)+3=3(2^{2e}-1)$. On the other hand the Fermat curve is non-singular and hence, by the Hasse-Weil bound,
$N\leq q^{2}+1+qs(s-1)=2^{2e}+2^{2e-1}-2^{\frac{3e-1}{2}}+1.$
Nevertheless, $2^{2e}+2^{2e-1}-2^{\frac{3e-1}{2}}+1\leq 3(2^{2e}-1)$ for $e\geq 3$. Therefore, this leads to a contradiction. Hence, it follows that there is at least an element $\alpha$ such that the corresponding line $\ell_{\alpha}$ is external to $\mathscr{F}^{2}_{s+1}$.
4.2 Intersection of $\mathscr{V}^{4}_{\varepsilon}$ with planes
In the following, we show that $\tau_{2}(\mathscr{V}^{4}_{\varepsilon})=0$ studying how planes meet $\mathscr{V}^{4}_{\varepsilon}$.
Lemma 4.6.
In $\mathrm{PG}(4,q^{2})$, there exists a plane that meets the variety $\mathscr{V}^{4}_{\varepsilon}$ in exactly one point.
Proof.
By the results of Subsection 4.1, there exists a line $\ell$ external to the Fermat curve $\mathscr{F}^{2}_{s+1}$ in the plane $\Pi_{0}:X_{0}=X_{4}=0$. In order to prove the statement, it is enough to show that the plane $\Pi=\langle P_{\infty},\ell\rangle$ meets $\mathscr{V}_{\varepsilon,\infty}$ in one point. By construction $P_{\infty}$ belongs to $\Pi\cap\mathscr{V}^{4}_{\varepsilon,\infty}$; suppose that there exists a point $Q\in\Pi\cap\mathscr{V}^{4}_{\varepsilon,\infty}$, $Q\neq P_{\infty}$. Since $\mathscr{V}^{4}_{\varepsilon,\infty}$ is a cone with vertex $P_{\infty}$ over $\mathscr{F}^{2}_{s+1}$ than there exists a point $R\in\mathscr{F}^{2}_{s+1}$ such that $Q$ belongs to the line $P_{\infty}R$. Then $P_{\infty}Q=P_{\infty}R\subseteq\Pi$ and this meets the line $\ell$ in the point $R$, getting a contradiction. ∎
Theorem 4.7.
Each plane in $\mathrm{PG}(4,q^{2})$ meets the variety $\mathscr{V}^{4}_{\varepsilon}$ in at least one point.
Proof.
Suppose there exists a plane $\Pi$ disjoint from the variety $\mathscr{V}^{4}_{\varepsilon}$. This plane must meet the hyperplane at infinity in a line. Consider the $q^{2}+1$ hyperplanes of $\mathrm{PG}(4,q^{2})$ containing $\Pi$. By Theorem 3.9 for $r=4$, each of these planes meets $\mathscr{V}^{4}_{\varepsilon}\cap\mathrm{AG}(4,q^{2})$ in $q^{5}-q^{3}+q^{2}$, or $q^{5}$, or $q^{5}+q^{2}$ affine points, Since the cardinality of $\mathscr{V}^{4}_{\varepsilon}$ is $q^{7}$ it follows that
$q^{7}=|\mathscr{V}^{4}_{\varepsilon}\cap\mathrm{AG}(4,q^{2})|\geq(q^{2}+1)(q^{5}-q^{3}+q^{2})>q^{7},$
a contradiction. ∎
Lemma 4.8.
Each line of $\mathrm{PG}(2,q^{2})$ meets the Fermat curve $\mathscr{F}^{2}_{s+1}$ in $0,1,2$ or $t$ points where $t=5$ if $e\equiv 1\pmod{4}$ and $e>1$ or $t=3$ if $e\equiv 3\pmod{4}$.
Proof.
We need to solve the following system in $X_{1},X_{2},X_{3}$:
$\begin{cases}X_{1}^{s+1}+X_{2}^{s+1}+X_{3}^{s+1}=0\\ \alpha X_{1}+\beta X_{2}+\gamma X_{3}=0.\end{cases}$ (31)
Since all the expressions are symmetric in $X_{1},X_{2},X_{3}$ we can assume without loss of generality $\gamma\neq 0$ i.e., equivalently, set $\gamma=1$. Substituting the second equation of System (31) in the first we obtain
$(1+\alpha^{s+1})X_{1}^{s+1}+\alpha^{s}\beta X_{1}^{s}X_{2}+\alpha\beta^{s}X_{1}X_{2}^{s}+(1+\beta^{s+1})X_{2}^{s+1}=0.$ (32)
The number of solutions of Equation (32) corresponds to the number of absolute points in $\mathrm{PG}(1,2^{2e})$ of a (possibly degenerate) $s$-sesquilinear form. By [14, Lemma 6] this corresponds to either $0,1,2$ or $t$ points where $t$ is the cardinality of a subline over the fixed subfield of $s$. In particular, for $e\equiv 1\pmod{4}$ we have $t=5$ and for $e\equiv 3\pmod{4}$ we have $t=3$. ∎
We can now specialize the results of Theorem 3.12 to the case $r=4$.
Proposition 4.9.
The linear code $\mathscr{C}^{4}_{\varepsilon}=\mathscr{C}(\mathscr{V}^{4}_{\varepsilon})$ for $q=2^{e}$ and $e>1$ an odd integer is a $5$-dimensional minimal code with the following parameters
  1. 1.
    If $e\equiv 1\pmod{4}$, then $\mathscr{C}^{4}_{\varepsilon}$ has length $|\mathscr{V}^{4}_{\varepsilon}|=q^{7}+q^{4}+q^{2}+1$ and weights
    $\{q^{7},q^{7}-q^{5}+q^{4}+q^{3}-q^{2},q^{7}-q^{5}+q^{4}+q^{2},\\ q^{7}-q^{5}+q^{4},q^{7}-q^{5}+q^{4}-q^{2},q^{7}-q^{5}+q^{4}-4q^{2}\};$
  2. 2.
    If $e=3$, then $\mathscr{C}^{4}_{\varepsilon}$ has length $|\mathscr{V}_{\varepsilon}^{4}|=q^{7}+q^{4}+2q^{3}+q^{2}+1$ and weights
    $\{q^{7},q^{7}-q^{5}+q^{4}+3q^{3}-q^{2},q^{7}-q^{5}+q^{4}+2q^{3}+q^{2},\\ q^{7}-q^{5}+q^{4}+2q^{3},q^{7}-q^{5}+q^{4}+2q^{3}-q^{2},q^{7}-q^{5}+q^{4}+2q^{3}-2q^{2}\}.$
Proof.
The result follows from Theorem 3.10, Theorem 3.9, Proposition 3.11 and Lemma 4.6. In order to compute the weights, it is enough to recall that
$|\Pi_{\infty}\cap\mathscr{V}_{\varepsilon}^{4}|=1+q^{2}|\mathscr{F}_{n}^{2}|$
and
$|\Pi\cap\mathscr{V}_{\varepsilon}^{4}|=|\Pi\cap\Pi_{\infty}\cap\mathscr{V}_{\varepsilon}^{4}|+|\Pi\cap\mathscr{H}^{4}_{\varepsilon}|-|\Pi\cap\Pi_{\infty}\cap\mathscr{H}_{\varepsilon}^{4}|$
for any hyperplane $\Pi$ distinct from $\Pi_{\infty}$.
Finally, by the results obtained above, we can state the following.
Proposition 4.10.
Consider the variety $\mathscr{V}^{4}_{\varepsilon}$. Then
$\tau_{1}(\mathscr{V}^{4}_{\varepsilon})=2,\quad\tau_{2}(\mathscr{V}^{4}_{\varepsilon})=2,\quad\tau_{3}(\mathscr{V}^{4}_{\varepsilon})=0.$ (33)
Proof.
The first cutting gap follows by Subsection 4.1. The second one follows by Lemma 4.6 and Theorem 4.7. Finally, by Theorem 3.9, we have the third cutting gap. ∎
Note that since there are lines external to $\mathscr{V}^{4}_{\varepsilon}$ we have $\tau_{1,0}(\mathscr{V}^{4}_{\varepsilon})=1\neq\tau_{1}(\mathscr{V}^{4}_{\varepsilon})$.